OpenStudy (aaronandyson):

1200/x- 40 - 1200/x = 20/60

OpenStudy (aaronandyson):

\[\frac{ 1200 }{ x - 40 } - \frac{ 1200 }{ x } = \frac{ 20 }{ 60 }\]

OpenStudy (aaronandyson):

@rvc

OpenStudy (aaronandyson):

@zepdrix

OpenStudy (zzr0ck3r):

multiply every term by \(60x(x-40)\) \[60x(x-40)(\frac{ 1200 }{ x - 40 }) - 60x(x-40)(\frac{ 1200 }{ x }) = 60x(x-40)(\frac{ 20 }{ 60 })\] \[60x\cancel{(x-40)}(\frac{ 1200 }{ \cancel{x - 40} }) - 60\cancel{x}(x-40)(\frac{ 1200 }{ \cancel{x} }) = \cancel{60}x(x-40)(\frac{ 20 }{ \cancel{60} })\] \[60x(1200)-60(x-40)1200=x(x-40)(20)\]

OpenStudy (zzr0ck3r):

can you get it from here?

OpenStudy (aaronandyson):

#justgotlost :(

OpenStudy (zzr0ck3r):

I am going to need more than that to help you

OpenStudy (aaronandyson):

???

OpenStudy (zzr0ck3r):

I typed a bunch of stuff and your response is a hashtag, which tells me exactly nothing about where you are confused. If you clarify then I might be able to help

OpenStudy (aaronandyson):

Look,I'm doing this chapter/topic after MONTHS , so I have no idea as to what to do. And Why did you take The LCM as 60x(x-40)?

OpenStudy (zzr0ck3r):

Because that is what it is.

OpenStudy (zzr0ck3r):

you dont have to find the least, but it will save you time often. you can ALWAYS just take the product of the denominators. i.e. given \[\frac{1}{a}+\frac{1}{b}=\frac{1}{c}\] Then we have \[abc\frac{1}{a}+abc\frac{1}{b}=abc\frac{1}{c}\] \[bc+ac=ab\]

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (aaronandyson):

Yes.

OpenStudy (zzr0ck3r):

If you have any questions, ask Galois (@rational). I got to sleep. Good luck

OpenStudy (aaronandyson):

@rational @Michele_Laino

OpenStudy (aaronandyson):

I'll write the same equation as \[\frac{ 1200 }{ x } - \frac{ 1200 }{ x - 40 } = \frac{ 1 }{ 3 }\] Can't I?

OpenStudy (michele_laino):

yes you can!

OpenStudy (aaronandyson):

So then \[84,000 = x^2 + 40x ?!?!?!\]

OpenStudy (michele_laino):

there is an error of sign, since your equation is: \[\frac{{1200}}{{x - 40}} - \frac{{1200}}{x} = \frac{1}{3}\]

OpenStudy (michele_laino):

now, the least common multiple is: 3x(x-40), so your equation is equivalent to this one: \[1200 \cdot 3x - 1200 \cdot 3\left( {x - 40} \right) = x\left( {x - 40} \right)\]

OpenStudy (aaronandyson):

x^2 + 40x - 84,000 = 0?

OpenStudy (michele_laino):

no, we have: \[{x^2} - 40x - 144000 = 0\]

OpenStudy (aaronandyson):

How??

OpenStudy (michele_laino):

3*1200*40= 144000

OpenStudy (aaronandyson):

Oh I got my mistake. What next how do I even factor this!?!

OpenStudy (michele_laino):

you have to solve that equation above, using this formula: \[x = \frac{{40 \pm \sqrt {{{40}^2} + 4 \times 144000} }}{2} = ...?\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (aaronandyson):

I'm suppose to factor it and not use the formula -_-

OpenStudy (michele_laino):

it is the sam, since if you want to factor a polynomial, in general,, you have to find the roots of that polynomial

OpenStudy (michele_laino):

same*

OpenStudy (michele_laino):

hint: \[\Large \sqrt {{{40}^2} + 4 \times 144000} = 760\]

OpenStudy (aaronandyson):

"I'm suppose to factor it and not use the formula -_-"

OpenStudy (michele_laino):

I understand, nevertheless, in order to factorize your polynomial, you have to compute the roots of your polynomial, first

OpenStudy (aaronandyson):

What?

OpenStudy (michele_laino):

performing the computation which I wrote above, namely: \[x = \frac{{40 \pm \sqrt {{{40}^2} + 4 \times 144000} }}{2} = \frac{{40 \pm 760}}{2} = ...?\]

OpenStudy (michele_laino):

you should get two values

OpenStudy (aaronandyson):

Can we factor it please?pleeeeeeeeeeeeease?

OpenStudy (michele_laino):

\[\begin{gathered} {x_1} = \frac{{40 + 760}}{2} = ...? \hfill \\ \hfill \\ {x_2} = \frac{{40 - 760}}{2} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please, we need to know those two values above, in order to make the factorization

OpenStudy (michele_laino):

here is your factorization: \[{x^2} - 40x - 144000 = \left( {x - {x_1}} \right)\left( {x - {x_2}} \right)\]

OpenStudy (aaronandyson):

400 -360

OpenStudy (michele_laino):

perfect! So your factorization is: \[{x^2} - 40x - 144000 = \left( {x + 360} \right)\left( {x - 400} \right)\]

OpenStudy (aaronandyson):

???????????????????????????????????????????????????????????????????????????????????????? I thought it would be - 360 and +400 Whats wrong ????????????????????????????????????????????????????????????????????????????????????????

OpenStudy (michele_laino):

you have to replace x_1 with -360, and x_2 with 400 into this formula: \[{x^2} - 40x - 144000 = \left( {x - {x_1}} \right)\left( {x - {x_2}} \right)\] what do you get?

OpenStudy (aaronandyson):

idk :(

OpenStudy (michele_laino):

here is more steps: \[\begin{gathered} \left( {x - {x_1}} \right)\left( {x - {x_2}} \right) = \left\{ {x - \left( { - 360} \right)} \right\}\left( {x - 400} \right) = \hfill \\ \hfill \\ = \left( {x + 360} \right)\left( {x - 400} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

here are*

OpenStudy (aaronandyson):

so v r look at 400 then?